The following equation can clearly be simplified: In{ lim{f(x)-- - (X-1] +$}} + (sinº q + cos?q) + (sinºq + cos?q) = Ë cos
Answer #1

Answer: Since p-value > \boldsymbol{\alpha} there is insufficient evidence that the proportion is different from 7%

Explanation:

Null Hypothesis, Ho: The proportion of Mathematicians who can solve the equation is 0.07

\boldsymbol{i.e. P = 0.07}

Alternative Hypothesis, Ho=a: The proportion of Mathematicians who can solve the equation is significantly different from 0.07

\boldsymbol{i.e. P \neq 0.07}

Test-Statistics: Under Ho,

  р Z = P P*(1-P) 2 N(0, 1) n

  \boldsymbol{\hat{p} =\frac{42}{500}=0.084}

Therefore,

  \boldsymbol{Z = \frac{0.084-0.07}{\sqrt{\frac{0.07*0.93}{500}}} = 1.227}

\boldsymbol{P-value = P[|Z| > 1.227] = 0.2198}